Two written questions that can be given to learners for classwork to test their understanding of the subtraction of fractions.

1
Expert's answer
2022-01-18T13:58:17-0500

This depend on the year a learner is.

The questions below are suitable for year 5 learners where they are tested on understanding of subtraction and addition of fractions with the same denominator and different denominator.


a) "\frac {1}{5}+ \frac {2}{5}"


b) "\frac {6}{8} -\frac {1}{4}"